OpenStudy (anonymous):

PLEASE HELP ME ****need this to be 100% right please***

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

c

OpenStudy (anonymous):

@anx @midhun.madhu1987 @Nurali

OpenStudy (michele_laino):

please, you have to use the subsequent inverse formula: \[C=\frac{ 5 }{9 }(F-32)\]

OpenStudy (michele_laino):

please, try to substitute your fahrenheit temperatures, and get the correspondent celsius temperatures

OpenStudy (michele_laino):

for example: \[t _{1}=\frac{ 5 }{ 9 }(49-17)\approx 10\]

OpenStudy (michele_laino):

Sorry, an error is occurred! I retry: \[t _{1}=\frac{ 5 }{ 9 }(49-32) \approx 10\]

OpenStudy (anonymous):

thanks

OpenStudy (michele_laino):

other temperature is: \[t _{2}=\frac{ 5 }{ 9 }(72-42) \approx 21\] so the difference is : 21-10=11